OpenStudy (anonymous):

can some one please help me ? pretty please What is the equation of the line that passes through (-2, 2) and (1, -4)?

OpenStudy (anonymous):

To calculate the equation in slope intercept form let x1=-2 y1=2 x2= 1 y2=-4 1. Find the slope of the points using slope formula slope(m) = Rise / Run slope(m) = (y2 - y1)/(x2 - x1) 2. Once you have m you can convert into your equation using this y – y1 = m(x – x1) Then rearrange and simplify into y = mx + b

OpenStudy (anonymous):

b will be your constants that are left over.. y and x are just the y and x from the equation, and m will be your slope

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

What is the length of AB?

OpenStudy (anonymous):

IM CLUE LESS IN MATH stuk in summerschool nd cant figure this out

OpenStudy (anonymous):

i have to finish this odysseyWare

OpenStudy (anonymous):

for that one you need pythagoras rule a^2 + b^2 = c^2 Note that the 10unit line is a right angle to the diagonal, and therefore it must be positioned at the distance of the radius ...  so we know 2 sides of the triangle let's say we know A and C and we want to know B if a+b=c then c-a=b so b^2 = c^2 - a^2 b^2 = 14.5^2 - 10^2 Once we have b^2 we can just take the square root to find b b = SQRT(b)

OpenStudy (anonymous):

once you know the 3rd side.. you just subtract that from 14.5 and you have AB

OpenStudy (anonymous):



OpenStudy (anonymous):

thats where the a^2 + b^2 = c^2 is coming from if you are wondering.